A Complete Bibliography of Publications in Science Studies

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

04 December 2018 Version 1.00

Title word cross-reference

1970-71 [FG73]. 1970/71 [Roc74]. 1970s [dSP71, FOC^{+}71].

Absurdum} [Mey74]. Academic [dSP71]. Academics [Rif72]. Accessibility [Ano71a, Ano71b, Ano71c, Ano72a, Ano72b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74b, Ano74c, Ano74d]. ad [Mey74]. Advancement [Ora71]. Advisory [Mul72]. Affair [Wil73]. after [SG73]. Agriculture [Sch72a]. Air [Eri72]. Allocations [Orl73]. American [Ezr71, FG73]. Amerikanische [Pfe71]. among [Rif72, Tür73]. Analysis [Bri72, Kra72]. Announcement [Gar73b, ME74]. Applicable [Bri72]. Applied [WG74]. Approach [Lon71]. Archaeology [Jah72]. Archives [Ano72e]. Array [Sch72b]. Arts [Hud73]. Aspects [Lew72, Wil73]. Association [Ora71]. Attitudes [Zin71, Tua71]. Australian [Hil74]. Authors [SR72]. Aversion [JLM73].

Back [Ano71d, Ano74e]. Background [Ora71]. Bank [Gar73b]. Behaviour [Gar73b]. Bibliographical [MO71]. Biology [You71]. Bleak [Gar73a]. Bloor [Mey74]. Book [Blo71, Blo74, Dea74, Eas74, Eri71, Haw73, Hol73, Jah72, Kea73, Pfe71, Pri71, Rav74, Sch72b, Tua71]. Bose [Nan72]. British [Ora71, San72, Wer71]. Burnet [JL72].

Canadian [Gar73a]. Careers [Zin71]. Case [GK72, Hol74, Law73, Mul73]. Catalogue [dB72]. Census [Roc74]. Challenge [SG73]. Chandra [Nan72]. Change [Ano72e, NHM74]. Chemistry [Zin71]. China [Dea74]. Choice [JLM73]. Civic [RS74]. Climate [Gar73a]. Comment [JM73]. Commentary [Rif72]. Commission [SG74a, Sko73]. Commitment [Zin71]. Communication [Gar73b]. Communications [WG74]. Comparative [Lon71]. Complex [Eri71]. Conceptual [Mul74, SR73]. Conditions [Wei73]. Conformity [Nan72]. Context [SR73]. Continuity [GT72]. Contradiction [Rif72]. Contrary [JM73]. Control [NHM74]. Controversy [GK72, Nel71]. Countries [Sag73]. Course [Eas73]. Criterion [Dol74]. Crystallography [Law73]. Cultural [Sha72]. Cybernetic [PT72]. Cybernetics [Hol74].

D [Orl73, FOC^{+}71]. Data [Gar73b, SR72]. David [Mey74]. Debate [War74, WP74]. Defence [Dol74, Eri72]. Defiance [Nan72]. Defines [MR73]. Delhi [SG74a]. Democratic [SR73]. Design [Eri72]. Development [Eri72, GK72, Law73, Lew72, Mul73, SR73]. Disarray [Sch72b]. Discovery [Dun74]. Displacement [Mul74]. Doctoral [Tür73]. Doctorate [Tür73]. Doubts [War71]. Durrant [Pri71].

Early [Ber72]. Edinburgh [Mor73]. Editorial [Ano71e, EM73, ME74]. Effort [Lew72]. Elite [Mul72]. Elitism [Ami74]. English [RS74]. Environment [Tür73]. Environmental [Nel71, Tua71]. Ethical [Sch72b]. Ethnomethodology [Mul73]. Europe [SR72]. Evaluation [Ber72]. Evolution’ [Wei73]. Evolutionary [You71]. Examination [GW74, Sha72].

Factor [JLM73]. Fertile [JM73]. Fertility [Hud73]. Findings [JM73]. Firm [San72]. Firmament [Pri71]. First [Eas73, Zin71, SG73]. First-Year [Eas73, Zin71]. Foucault [Jah72]. Front [Ano71f, Ano71g, Ano71h, Ano72f, Ano72g, Ano72h, Ano72i, Ano73e, Ano73f, Ano73g, Ano73h, Ano74f, Ano74g, Ano74h, Ano74i, Sch72b]. Funding [dSP71].

Gap [Zin71]. Genius [Woo74]. Geoffrey [Pri71]. German [SR73]. Germanies [SR72]. Goals [FOC^{+}71]. Government [Hut71, Lon71]. Graduate [FG73]. Graphing [SG74b]. Great [Pri71]. Growth [MO71]. Guide [MO71].

Health [Mul72]. Henry [MR73]. High [Pri71]. Hindsight [Kre71].

Historiography [Wei74a, Wei74b]. History [Eas73, FG73, Mir72, Wei74a, Wei74b]. Human [Jah72]. Hundred [Dea74].

Idea [Dun74]. Identifying [SG74b]. Identity [Nan72]. Ideology [You71]. II [GSSD74, Wei74b]. Imaginations [JM73]. Impact [NHM74]. Implications [PT72, Roc74]. Index [Ami74]. Indicator [SR72]. Indifference [Dun74]. Industrial [Bar71, Eri71]. Industry [Gar73a, San72]. Influence [Hil74, Hut71]. Informal [WG74]. Information [Ano71i, Ano72j, Ano73i, Ano74j]. Ingenious [MR73]. Innovation [Hol74]. Institution [Ber72]. Institutional [Ano72e, SR73]. Intellectual [Kra72]. Interest [Haw73]. International [SG73, SG74a, Sko73]. Internationale [Sal71]. Introduction [Eas73]. Isolation [Dun74].

J [Pri71]. Jacques [Rav74]. Jagadis [Nan72]. Jean [Rav74]. Jean-Jacques [Rav74]. Joseph [MR73]. Journal [SR72].

Knowledge [Col74, Dol71, Sch72a, Blo71, Blo74, Haw73]. Kuhnian [Bri72].

Laboratories [Hut71]. Lessons [Wei74b]. Liberation [Eas74]. Literature [GW74]. Literatures [SG74b, GSSD74]. Location [Eri72]. Loss [Sil74, Woo74]. Loyalties [SG73].

Machine [Gar73b]. Machine-Readable [Gar73b]. Macro [GSSD74]. Macro- [GSSD74]. Making [Bar71]. Manpower [RS74, SR72]. Mathematical [Kra72]. Matter [Ano71d, Ano71f, Ano71g, Ano71h, Ano72f, Ano72g, Ano72h, Ano72i, Ano73e, Ano73f, Ano73g, Ano73h, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i]. Meadows [Pri71]. Mechanics [MR73]. Medico [Sch72b]. Meeting [SG74a]. Merton [Kea73]. Methodological [SR72]. Methodology [Wei74b]. Michel [Jah72]. Microstructure [GSSD74]. Mid [Mor73]. Mid-Victorian [Mor73]. Migration [Mul74]. Military [Eri71]. Millenarianism [JL72]. Misparadigmatic [Lip74]. Mission [Ami74]. Mystification [Blo74].

Natural [Dol71]. Nature [Eas74, Wei74a]. Needs [Eas74]. Networks [Col74]. Norm [Woj74]. Normal [War74, WP73]. Norms [Woj74]. North [FG73]. Note [EM73].

Objective [Blo74, Mey74]. Objectivity [Dol74]. One [Dea74]. Order [Jah72]. Organizations [SG73]. Organized [Hut71]. Origin [dB72]. Outsider [Wer71]. Ovonic [GK72].

Paper [Mul74]. Papers [dB72]. Paradigms [Blo71, WP73]. Parent [Woo74]. Parental [Sil74]. Patronage [Mor73]. Peter [Pfe71]. Philosophical [Sha72]. Physicists [Tür73]. Place [Pal71, War71]. Point [Sag73]. Points [MO71]. Policy [Gar73a, Kre71, Kre73, SG74a, Sko73, SR73, WG74]. Political [Ezr71, JL72, SR73, Wil73]. Politics [Dea74, Wer71, Rav74]. Popper [Blo74]. Post [Tür73]. Post-Doctorate [Tür73]. Postparadigmatic [Lip74]. Pottery [Sha72]. Prefatory [Mul74]. Preparadigmatic [Lip74]. Principles [dSP71]. Problem [Eri72]. Problematic [Wei74a]. Problems [Mir72]. Productivity [GT72, Tür73]. Professional [PT72]. Projecting [dSP71]. Prospects [Gar73a, Mir72]. Protein [Law73]. Provincial [Ora71, Sha72]. Psychology [Bri72, Lip74, Pal71, War71, WP73, Wei74a, Wei74b]. Public [Mul72].

Quantitative [GW74]. Questioning [Hil74]. Questions [Dea74].

R [Orl73, FOC^{+}71]. Radar [Eri72]. Radio [Eri72]. Radio-Location [Eri72]. Rationality [Woj74]. Ray [Law73]. Re [Ber72]. Re-Evaluation [Ber72]. Readable [Gar73b]. Real [Kre71]. Reductio [Mey74]. Reflections [Dea74]. Rejoinder [WP74]. Republic [SR73]. Research [Bar71, Lew72, Mul71, San72, Wei74b]. Reservations [War71]. Resistance [Dun74]. Resources [Ezr71]. Retrieval [Wei74a, Wei74b]. Review [Blo71, Blo74, Dea74, Eas74, Eri71, Haw73, Hol73, Jah72, Kea73, Pfe71, Pri71, Rav74, Sch72b, Tua71]. Revisited [Kea73]. Revolution [Pal71, War71]. Rothschild [Wil73]. Royal [Ber72, dB72]. Russian [Hol73].

Sacred [JL72]. Salomon [Rav74]. Scholarly [War74]. Scholarship [WP74]. School [JLM73]. Science [Dol71, Eas73, Ezr71, FG73, Gar73a, GW74, Hil74, Hol74, Kra72, Kre71, Kre73, Law73, JLM73, Lon71, MO71, Mir72, MR73, Mor73, Mul74, Mul73, Nan72, dSP71, Roc74, Sag73, Sal71, SG74a, Sha72, Sko73, SR73, GSSD74, War74, WP73, WG74, Zin71, Ora71, Dea74, Hol73, Rav74]. Sciences [Hud73, Jah72]. Scientific [Ami74, Blo71, Col74, Dol74, Dun74, GK72, Haw73, Pal71, Roc74, RS74, SG73, SG74b, SR72, GSSD74, War71, Wei74b, Wei73, Woo74, dB72]. Scientist [Ami74]. Scientists [GT72, Gar73b, Hil74, Hut71, Nel71, Sil74, Wer71]. Seminar [Eas73]. Service [Mul72]. Set [Col74]. Sexuality [JM73]. Smithsonian [dB72]. Social [Dol74, Eas73, Hil74, Kra72, Mul73, PT72]. Society [dB72, Sha72, Hol73]. Sociological [Mul71]. Sociology [Dol71]. Some [Lew72, Mul71, Wei74b, Wil73]. Soviet [Hol73, Eri72, Hol74, Lew72]. Specialization [JM73]. Specialties [Law73, Mul73, SG74b]. Staff [Zin71]. Standards [War74, WP74]. State [PT72]. States [Orl73]. Statistics [MO71]. Structure [Mul72, SG74b, GSSD74]. Structures [Kra72]. Students [Eas73, FG73, Zin71]. Studies [Eas73, Mir72, RS74, SG74a, SR73, Sko73]. Study [GK72, GW74, Hil74, SR72]. Subject [JM73]. Subjects [JLM73].

Suggestions [Mul71]. Survey [FG73]. Switches [GK72]. Symposium [Wei73]. System [Hil74, Pri71].

Tacit [Col74]. Taking [Pal71, War71]. TEA [Col74]. Teaching [JLM73]. Technological [NHM74]. Technologists [Gar73b]. Technology [Sag73, WG74]. Theorists [MR73]. Theory [JL72, Kre73]. Things [Jah72]. Thousand [Dea74]. Training [Tür73]. Transnational [SG73]. Truth [Mey74]. Turkish [Tür73]. Two [SR72, Blo71].

U.S. [Mul72]. Underdeveloped [Sag73]. Underdevelopment [Sag73]. Union [Hol74, Lew72]. United [Orl73]. Universities [RS74]. University [Mor73]. Uses [Sha72]. Using [SR72]. USSR [Mir72].

Venezuela [Roc74]. Victorian [Mor73]. View [Sag73]. Volume [Ano71i, Ano72j, Ano73i, Ano74j].

War [SG73]. Warren [WP74]. Weingart [Pfe71]. Welfare [Sch72a]. Who [Eas74]. Widening [Zin71]. Wintry [Gar73a]. Wissenschaftslobby [Pfe71]. Wordsworth [Pri71]. Work [PT72]. Workshop [Rif72]. World [Kre71, SG73].

X [Law73]. X-Ray [Law73].

Year [Eas73, Zin71]. Years [Ber72, GT72].

References

Amick:1974:ISE

[Ami74] Daniel J. Amick. An index of scientific elitism and the scientist’s mission. Science Studies, 4(1):1–16, January 1974. CODEN ????

         ISSN 0036-8539. URL https://www.jstor.org/stable/284535.

Anonymous:1971:Aa

[Ano71a] Anonymous. Accessibility. Science Studies, 1(1):??, January 1971. CODEN ???? ISSN 0036-8539.

Anonymous:1971:Ab

[Ano71b] Anonymous. Accessibility. Science Studies, 1(2):??, April 1971. CO-DEN ???? ISSN 0036-8539.

Anonymous:1971:Ac

[Ano71c] Anonymous. Accessibility. Science Studies, 1(3/4):??, October 1971. CODEN ???? ISSN 0036-8539.

Anonymous:1971:BM

[Ano71d] Anonymous. Back matter. Science Studies, 1(2):243–244, April 1971. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/370211.

Anonymous:1971:E

[Ano71e] Anonymous. Editorial. Science Studies, 1(1):1–2, January 1971. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/370193.

Anonymous:1971:FMa

[Ano71f] Anonymous. Front matter. Science Studies, 1(1):i–x, January 1971. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/370192.

Anonymous:1971:FMb

[Ano71g] Anonymous. Front matter. Science Studies, 1(2):??, April 1971. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/370201.

Anonymous:1971:FMc

[Ano71h] Anonymous. Front matter. Science Studies, 1(3/4):??, October 1971. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/370212.

Anonymous:1971:VI

[Ano71i] Anonymous. Volume information. Science Studies, 1(1):iii–vii, January 1971. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/370191.

Anonymous:1972:Aa

[Ano72a] Anonymous. Accessibility. Science Studies, 2(1):??, January 1972. CODEN ???? ISSN 0036-8539.

Anonymous:1972:Ab

[Ano72b] Anonymous. Accessibility. Science Studies, 2(2):??, April 1972. CO-DEN ???? ISSN 0036-8539.

Anonymous:1972:Ac

[Ano72c] Anonymous. Accessibility. Science Studies, 2(3):??, July 1972. CO-DEN ???? ISSN 0036-8539.

Anonymous:1972:Ad

[Ano72d] Anonymous. Accessibility. Science Studies, 2(4):??, October 1972. CODEN ???? ISSN 0036-8539.

Anonymous:1972:AIC

[Ano72e] Anonymous. Archives of institutional change. Science Studies, 2 (1):102–103, January 1972. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/284492.

Anonymous:1972:FMa

[Ano72f] Anonymous. Front matter. Science Studies, 2(1):1, January 1972. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284487.

Anonymous:1972:FMb

[Ano72g] Anonymous. Front matter. Science Studies, 2(2):??, April 1972. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284510.

Anonymous:1972:FMc

[Ano72h] Anonymous. Front matter. Science Studies, 2(3):??, July 1972. CO-

DEN ???? ISSN 0036-8539. URL https://www.jstor.org/stable/
284515.

Anonymous:1972:FMd

[Ano72i] Anonymous. Front matter. Science Studies, 2(4):??, October 1972. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284501.

Anonymous:1972:VI

[Ano72j] Anonymous. Volume information. Science Studies, 2(4):i–v, October 1972. CODEN ???? ISSN 0036-8539. URL https://www.jstor.

org/stable/284500.

Anonymous:1973:Aa

[Ano73a] Anonymous. Accessibility. Science Studies, 3(1):??, January 1973. CODEN ???? ISSN 0036-8539.

Anonymous:1973:Ab

[Ano73b] Anonymous. Accessibility. Science Studies, 3(2):??, April 1973. CO-DEN ???? ISSN 0036-8539.

Anonymous:1973:Ac

[Ano73c] Anonymous. Accessibility. Science Studies, 3(3):??, July 1973. CO-

DEN ???? ISSN 0036-8539.

Anonymous:1973:Ad

[Ano73d] Anonymous. Accessibility. Science Studies, 3(4):??, October 1973.

CODEN ???? ISSN 0036-8539.

Anonymous:1973:FMa

[Ano73e] Anonymous. Front matter. Science Studies, 3(1):??, January 1973. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284460.

Anonymous:1973:FMb

[Ano73f] Anonymous. Front matter. Science Studies, 3(2):??, April 1973. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284480.

Anonymous:1973:FMc

[Ano73g] Anonymous. Front matter. Science Studies, 3(3):??, July 1973. CO-

DEN ???? ISSN 0036-8539. URL https://www.jstor.org/stable/
284493.

Anonymous:1973:FMd

[Ano73h] Anonymous. Front matter. Science Studies, 3(4):??, October 1973. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284529.

Anonymous:1973:VI

[Ano73i] Anonymous. Volume information. Science Studies, 3(4):i–vi, October 1973. CODEN ???? ISSN 0036-8539. URL https://www.jstor.

org/stable/284528.

Anonymous:1974:Aa

[Ano74a] Anonymous. Accessibility. Science Studies, 4(1):??, January 1974.

CODEN ???? ISSN 0036-8539.

Anonymous:1974:Ab

[Ano74b] Anonymous. Accessibility. Science Studies, 4(2):??, April 1974. CO-

DEN ???? ISSN 0036-8539.

Anonymous:1974:Ac

[Ano74c] Anonymous. Accessibility. Science Studies, 4(3):??, July 1974. CO-DEN ???? ISSN 0036-8539.

Anonymous:1974:Ad

[Ano74d] Anonymous. Accessibility. Science Studies, 4(4):??, October 1974. CODEN ???? ISSN 0036-8539.

Anonymous:1974:BM

[Ano74e] Anonymous. Back matter. Science Studies, 4(2):201, April 1974. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284479.

Anonymous:1974:FMa

[Ano74f] Anonymous. Front matter. Science Studies, 4(1):??, January 1974. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284534.

Anonymous:1974:FMb

[Ano74g] Anonymous. Front matter. Science Studies, 4(2):??, April 1974. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284470.

Anonymous:1974:FMc

[Ano74h] Anonymous. Front matter. Science Studies, 4(3):??, July 1974. CO-

DEN ???? ISSN 0036-8539. URL https://www.jstor.org/stable/
284520.

Anonymous:1974:FMd

[Ano74i] Anonymous. Front matter. Science Studies, 4(4):338, October 1974. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284544.

Anonymous:1974:VI

[Ano74j] Anonymous. Volume information. Science Studies, 4(4):421–422, October 1974. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284543.

Barnes:1971:MIR

[Bar71] S. B. Barnes. Making out in industrial research. Science Studies, 1(2):157–175, April 1971. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/370204.

Berman:1972:EYR

[Ber72] Morris Berman. The early years of the Royal Institution 1799– 1810: A re-evaluation. Science Studies, 2(3):205–240, July 1972.

CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284516.

Bloor:1971:BRT

[Blo71] David Bloor. Book review: Two Paradigms for Scientific Knowledge? Science Studies, 1(1):101–115, January 1971. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/370200.

Bloor:1974:BRP

[Blo74] David Bloor. Book review: Popper’s Mystification of Objective Knowledge. Science Studies, 4(1):65–76, January 1974. CODEN ????

ISSN 0036-8539. URL https://www.jstor.org/stable/284538.

Briskman:1972:KAA

[Bri72] L. B. Briskman. Is a Kuhnian analysis applicable to psychology? Science Studies, 2(1):87–97, January 1972. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/284490.

Collins:1974:TST

[Col74] H. M. Collins. The TEA set: Tacit knowledge and scientific networks. Science Studies, 4(2):165–185, April 1974. CODEN ???? ISSN 0036-

8539. URL https://www.jstor.org/stable/284473.

deBBeaver:1972:SOR

[dB72] Donald deB. Beaver. The Smithsonian origin of the Royal Society Catalogue of Scientific Papers. Science Studies, 2(4):385–393, October 1972. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284508.

Dean:1974:BRS

[Dea74] Genevieve Dean. Book review: Science and Politics in China: Reflections on One Hundred Thousand Questions. Science Studies, 4 (1):93–96, January 1974. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/284540.

Dolby:1971:SKN

[Dol71] R. G. A. Dolby. Sociology of knowledge in natural science. Science Studies, 1(1):3–21, January 1971. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/370194.

Dolby:1974:DSC

[Dol74] R. G. A. Dolby. In defence of a social criterion of scientific objectivity. Science Studies, 4(2):187–190, April 1974. CODEN ???? ISSN 0036-

8539. URL https://www.jstor.org/stable/284474.

Price:1971:PPF

[dSP71] Derek J. de Solla Price. Principles for projecting funding of academic science in the 1970s. Science Studies, 1(1):85–94, January 1971. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/370198.

Duncan:1974:ISD

[Dun74] Simon S. Duncan. The isolation of scientific discovery: Indifference and resistance to a new idea. Science Studies, 4(2):109–134, April 1974. CODEN ???? ISSN 0036-8539. URL https://www.jstor.

org/stable/284471.

Easlea:1973:IHS

[Eas73] Brian Easlea. An introduction to the history and social studies of science: A seminar course for first-year science students. Science Studies, 3(2):185–209, April 1973. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/284486.

Easlea:1974:BRW

[Eas74] Brian Easlea. Book review: Who Needs the Liberation of Nature? Science Studies, 4(1):77–92, January 1974. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/284539.

Edge:1973:EN

[EM73] David Edge and Roy MacLeod. Editorial note. Science Studies, 3 (1):1, January 1973. CODEN ???? ISSN 0036-8539. URL https:/

/www.jstor.org/stable/284461.

Erickson:1971:BRM

[Eri71] J. Erickson. Book review: The Military–Industrial Complex. Science Studies, 1(2):225–233, April 1971. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/370208.

Erickson:1972:RLA

[Eri72] John Erickson. Radio-location and the air defence problem: The design and development of Soviet radar 1934–40. Science Studies, 2 (3):241–263, July 1972. CODEN ???? ISSN 0036-8539. URL https:

//www.jstor.org/stable/284517.

Ezrahi:1971:PRA

[Ezr71] Yaron Ezrahi. The political resources of American science. Science Studies, 1(2):117–133, April 1971. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/370202.

French:1973:SNA

[FG73] Richard French and Michael Gross. A survey of North American graduate students in the history of science 1970-71. Science Studies, 3(2):161–171, April 1973. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/284483.

Sinclair:1971:GRD

[FOC^{+}71] C. Freeman, C. H. G. Oldham, C. M. Cooper, T. C. Sinclair, and B. G. Achilladelis. The goals of R & D in the 1970s. Science Studies, 1(3/4):357–406, October 1971. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/370218.

Gardiner:1973:WPC

[Gar73a] J. Paul Gardiner. Wintry prospects for Canadian industry and science policy, in a bleak climate. Science Studies, 3(2):181–183, April 1973. CODEN ???? ISSN 0036-8539. URL https://www.jstor.

org/stable/284485.

Garvey:1973:AMR

[Gar73b] William D. Garvey. Announcement: A machine-readable data bank on the communication behaviour of scientists and technologists. Science Studies, 3(1):91–92, January 1973. CODEN ???? ISSN 0036-

8539. URL https://www.jstor.org/stable/284469.

Gibbons:1972:DOS

[GK72] Michael Gibbons and Philip King. The development of Ovonic switches: A case study of a scientific controversy. Science Studies, 2(4):295–309, October 1972. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/284502.

Stonehill:1974:SSL

[GSSD74] Belver C. Griffith, Henry G. Small, Judith A. Stonehill, and Sandra Dey. The structure of scientific literatures II: Toward a macro- and microstructure for science. Science Studies, 4(4):339–365, October 1974. CODEN ???? ISSN 0036-8539. URL https://www.jstor.

org/stable/284546.

Garvey:1972:CPS

[GT72] William D. Garvey and Kazuo Tomita. Continuity of productivity by scientists in the years 1968–71. Science Studies, 2(4):379–383, October 1972. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284507.

Gilbert:1974:QSS

[GW74] G. Nigel Gilbert and Steve Woolgar. The quantitative study of science: An examination of the literature. Science Studies, 4(3):279– 294, July 1974. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284526.

Hawthorn:1973:BRI

[Haw73] Geoffrey Hawthorn. Book review: The Interest of Scientific Knowledge. Science Studies, 3(1):78–87, January 1973. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/284467.

Hill:1974:QIS

[Hil74] Stephen C. Hill. Questioning the influence of a ‘social system of science’: A study of Australian scientists. Science Studies, 4(2): 135–163, April 1974. CODEN ???? ISSN 0036-8539. URL https:/

/www.jstor.org/stable/284472.

Holloway:1973:BRS

[Hol73] David Holloway. Book review: Science in Russian and Soviet Society. Science Studies, 3(1):61–72, January 1973. CODEN ???? ISSN 0036-

8539. URL https://www.jstor.org/stable/284465.

Holloway:1974:ISC

[Hol74] David Holloway. Innovation in science — the case of cybernetics in the Soviet Union. Science Studies, 4(4):299–337, October 1974. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284545.

Hudson:1973:FAS

[Hud73] Liam Hudson. Fertility in the arts and sciences. Science Studies, 3 (3):305–310, July 1973. CODEN ???? ISSN 0036-8539. URL https:

//www.jstor.org/stable/284497.

Hutchinson:1971:GLI

[Hut71] Eric Hutchinson. Government laboratories and the influence of organized scientists. Science Studies, 1(3/4):331–356, October 1971. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/370217.

Jahoda:1972:BRO

[Jah72] Marie Jahoda. Book review: The Order of Things: An Archaeology of the Human Sciences by Michel Foucault. Science Studies, 2(1): 99–101, January 1972. CODEN ???? ISSN 0036-8539. URL https:

//www.jstor.org/stable/284491.

Jacob:1972:PMB

[JL72] M. C. Jacob and W. A. Lockwood. Political millenarianism and Burnet’s sacred theory. Science Studies, 2(3):265–279, July 1972. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284518.

Littlejohn:1973:AST

[JLM73] C. L. Jones, G. M. Littlejohn, and A. F. McPherson. Aversion to school teaching as a factor in the choice of science subjects. Science Studies, 3(2):173–179, April 1973. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/284484.

Jones:1973:FIC

[JM73] C. L. Jones and A. F. McPherson. Fertile imaginations and contrary findings: A comment on subject specialization and sexuality. Science Studies, 3(4):389–391, October 1973. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/284532.

Kearney:1973:BRM

[Kea73] H. F. Kearney. Book review: Merton Revisited. Science Studies, 3(1):72–78, January 1973. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/284466.

Krauze:1972:SIS

[Kra72] Tadeusz K. Krauze. Social and intellectual structures of science — a mathematical analysis. Science Studies, 2(4):369–378, October 1972. CODEN ???? ISSN 0036-8539. URL https://www.jstor.

org/stable/284506.

Kreilkamp:1971:HRW

[Kre71] Karl Kreilkamp. Hindsight and the real world of science policy. Science Studies, 1(1):43–66, January 1971. CODEN ???? ISSN 0036-

8539. URL https://www.jstor.org/stable/370196.

Kreilkamp:1973:TTS

[Kre73] Karl Kreilkamp. Towards a theory of science policy. Science Studies, 3(1):3–29, January 1973. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/284462.

Law:1973:DSS

[Law73] John Law. The development of specialties in science: The case of X-ray protein crystallography. Science Studies, 3(3):275–303, July 1973. CODEN ???? ISSN 0036-8539. URL https://www.jstor.

org/stable/284496.

Lewis:1972:SAR

[Lew72] Robert A. Lewis. Some aspects of the research and development effort of the Soviet Union, 1924–35. Science Studies, 2(2):153–179, April 1972. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284512.

Lipsey:1974:PPP

[Lip74] Mark W. Lipsey. Psychology: Preparadigmatic, postparadigmatic, or misparadigmatic? Science Studies, 4(4):406–410, October 1974. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284549.

Long:1971:GSC

[Lon71] T. Dixon Long. The government of science: A comparative approach. Science Studies, 1(3/4):263–286, October 1971. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/370214.

MacLeod:1974:EA

[ME74] Roy MacLeod and David Edge. Editorial announcement. Science Studies, 4(3):203, July 1974. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/284521.

Meynell:1974:DBR

[Mey74] Hugo Meynell. David Bloor’s Reductio ad Absurdum of ‘objective’ truth. Science Studies, 4(2):190–193, April 1974. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/284475.

Mirsky:1972:SSU

[Mir72] E. M. Mirsky. Science studies in the USSR (history, problems, prospects). Science Studies, 2(3):281–294, July 1972. CODEN ????

ISSN 0036-8539. URL https://www.jstor.org/stable/284519.

Meadows:1971:BSG

[MO71] A. J. Meadows and J. G. O’Connor. Bibliographical statistics as a guide to growth points in science. Science Studies, 1(1):95–99, January 1971. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/370199.

Morrell:1973:PMV

[Mor73] J. B. Morrell. The patronage of mid-Victorian science in the University of Edinburgh. Science Studies, 3(4):353–388, October 1973. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284531.

Molella:1973:TIM

[MR73] Arthur P. Molella and Nathan Reingold. Theorists and ingenious mechanics: Joseph Henry defines science. Science Studies, 3(4):323– 351, October 1973. CODEN ???? ISSN 0036-8539. URL https://

www.jstor.org/stable/284530.

Mulkay:1971:SSS

[Mul71] Michael Mulkay. Some suggestions for sociological research. Science Studies, 1(2):207–213, April 1971. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/370206.

Mullins:1972:SEA

[Mul72] Nicholas C. Mullins. The structure of an elite: The advisory structure of the U.S. Public Health Service. Science Studies, 2(1):3–29, January 1972. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284488.

Mullins:1973:DSS

[Mul73] Nicholas C. Mullins. The development of specialties in social science: The case of ethnomethodology. Science Studies, 3(3):245–273, July 1973. CODEN ???? ISSN 0036-8539. URL https://www.jstor.

org/stable/284495.

Mulkay:1974:CDM

[Mul74] Michael Mulkay. Conceptual displacement and migration in science: A prefatory paper. Science Studies, 4(3):205–234, July 1974. CO-

DEN ???? ISSN 0036-8539. URL https://www.jstor.org/stable/
284522.

Nandy:1972:DCS

[Nan72] Ashis Nandy. Defiance and conformity in science: The identity of Jagadis Chandra Bose. Science Studies, 2(1):31–85, January 1972. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284489.

Nelkin:1971:SEC

[Nel71] Dorothy Nelkin. Scientists in an environmental controversy. Science Studies, 1(3/4):245–261, October 1971. CODEN ???? ISSN 0036-

8539. URL https://www.jstor.org/stable/370213.

Hershey:1974:ICT

[NHM74] Dorothy Nelkin, Cary Hershey, and Dennis Mueller. The impact and control of technological change. Science Studies, 4(1):97–103, January 1974. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284541.

Orange:1971:BAA

[Ora71] A. D. Orange. The British Association for the Advancement of Science: The provincial background. Science Studies, 1(3/4):315–329, October 1971. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/370216.

Orlans:1973:DRA

[Orl73] Harold Orlans. ‘D & R’ allocations in the United States. Science Studies, 3(2):119–159, April 1973. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/284482.

Palermo:1971:SRT

[Pal71] David S. Palermo. Is a scientific revolution taking place in psychology? Science Studies, 1(2):135–155, April 1971. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/370203.

Pfetsch:1971:BRA

[Pfe71] Frank Pfetsch. Book review: Die Amerikanische Wissenschaftslobby by Peter Weingart. Science Studies, 1(2):235–237, April 1971. CO-

DEN ???? ISSN 0036-8539. URL https://www.jstor.org/stable/
370209.

Prickett:1971:BRW

[Pri71] Stephen Prickett. Book review: Wordsworth and the Great System by Geoffrey Durrant; The High Firmament by A. J. Meadows. Science Studies, 1(2):237–241, April 1971. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/370210.

Pearman:1972:PIS

[PT72] J. R. Pearman and Winnifred M. Thomet. Professional implications for social work in a cybernetic state. Science Studies, 2(2):181–190, April 1972. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284513.

Ravetz:1974:BRS

[Rav74] J. R. Ravetz. Book review: Science and Politics by Jean-Jacques Salomon. Science Studies, 4(3):295–297, July 1974. CODEN ????

ISSN 0036-8539. URL https://www.jstor.org/stable/284527.

Rifkin:1972:CAA

[Rif72] Susan B. Rifkin. On ‘contradictions’ among academics (a commentary on a workshop). Science Studies, 2(4):395–399, October 1972. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284509.

Roche:1974:SVI

[Roc74] Marcel Roche. Science in Venezuela: Implications of the scientific census of 1970/71. Science Studies, 4(4):397–405, October 1974. CODEN ???? ISSN 0036-8539. URL https://www.jstor.org/

stable/284548.

Roderick:1974:SSS

[RS74] Gordon Roderick and Michael Stephens. Scientific studies and scientific manpower in the English civic universities 1870–1914. Science Studies, 4(1):41–63, January 1974. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/284537.

Sagasti:1973:UST

[Sag73] Francisco R. Sagasti. Underdevelopment, science and technology: The point of view of the underdeveloped countries. Science Studies, 3(1):47–59, January 1973. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/284464.

Salomon:1971:IS

[Sal71] Jean-Jacques Salomon. The internationale of science. Science Studies, 1(1):23–42, January 1971. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/370195.

Sanderson:1972:RFB

[San72] Michael Sanderson. Research and the firm in British industry, 1919– 39. Science Studies, 2(2):107–151, April 1972. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/284511.

Schultz:1972:KAW

[Sch72a] Theodore W. Schultz. Knowledge, agriculture and welfare. Science Studies, 2(4):361–368, October 1972. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/284505.

Schurr:1972:BRA

[Sch72b] George M. Schurr. Book review: Array and Disarray on the Medico– Ethical Front. Science Studies, 2(2):191–202, April 1972. CODEN

???? ISSN 0036-8539. URL https://www.jstor.org/stable/
284514.

Schroeder-Gudehus:1973:CTL

[SG73] Brigitte Schroeder-Gudehus. Challenge to transnational loyalties: International scientific organizations after the First World War. Science Studies, 3(2):93–118, April 1973. CODEN ???? ISSN 0036-

8539. URL https://www.jstor.org/stable/284481.

Schroeder-Gudehus:1974:ICS

[SG74a] Brigitte Schroeder-Gudehus. International Commission for Science Policy Studies: New Delhi Meeting. Science Studies, 4(1):105–107, January 1974. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284542.

Small:1974:SSL

[SG74b] Henry Small and Belver C. Griffith. The structure of scientific literatures I: Identifying and graphing specialties. Science Studies, 4(1): 17–40, January 1974. CODEN ???? ISSN 0036-8539. URL https:

//www.jstor.org/stable/284536.

Shapin:1972:PPS

[Sha72] Steven A. Shapin. The Pottery Philosophical Society, 1819–1835: An examination of the cultural uses of provincial science. Science Studies, 2(4):311–336, October 1972. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/284503.

Silverman:1974:PLS

[Sil74] S. M. Silverman. Parental loss and scientists. Science Studies, 4(3): 259–264, July 1974. CODEN ???? ISSN 0036-8539. URL https://

www.jstor.org/stable/284524.

Skolnikoff:1973:ICS

[Sko73] Eugene B. Skolnikoff. International Commission for Science Policy Studies. Science Studies, 3(1):89–90, January 1973. CODEN ????

ISSN 0036-8539. URL https://www.jstor.org/stable/284468.

Spiegel-Rosing:1972:JAI

[SR72] I. S. Spiegel-Rosing. Journal authors as an indicator of scientific manpower; a methodological study using data for the two Germanies and Europe. Science Studies, 2(4):337–359, October 1972. CODEN ????

ISSN 0036-8539. URL https://www.jstor.org/stable/284504.

Spiegel-Rosing:1973:SPS

[SR73] I. S. Spiegel-Rösing. Science policy studies in a political context: The conceptual and institutional development of science policy studies in the German Democratic Republic. Science Studies, 3(4):393–413, October 1973. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284533.

Tuan:1971:BRE

[Tua71] Yi-Fu Tuan. Book review: Environmental Attitudes. Science Studies, 1(2):215–224, April 1971. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/370207.

Turkeli:1973:DTE

[Tür73] Arif Türkeli. The doctoral training environment and post-doctorate productivity among Turkish physicists. Science Studies, 3(3):311– 318, July 1973. CODEN ???? ISSN 0036-8539. URL https://www.

jstor.org/stable/284498.

Warren:1971:SRT

[War71] Neil Warren. Is a scientific revolution taking place in psychology? — Doubts and reservations. Science Studies, 1(3/4):407–413, October 1971. CODEN ???? ISSN 0036-8539. URL https://www.jstor.

org/stable/370219.

Warren:1974:NSN

[War74] Neil Warren. Normal science and the normal standards of scholarly debate. Science Studies, 4(2):195–197, April 1974. CODEN ????

ISSN 0036-8539. URL https://www.jstor.org/stable/284477.

Weingart:1973:SCS

[Wei73] Peter Weingart. Symposium on ‘Conditions of Scientific Evolution’. Science Studies, 3(3):319–321, July 1973. CODEN ???? ISSN 0036-

8539. URL https://www.jstor.org/stable/284499.

Weimer:1974:HPRa

[Wei74a] Walter B. Weimer. The history of psychology and its retrieval from historiography: I. The problematic nature of history. Science Studies, 4(3):235–258, July 1974. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/284523.

Weimer:1974:HPRb

[Wei74b] Walter B. Weimer. The history of psychology and its retrieval from historiography II: Some lessons for the methodology of scientific research. Science Studies, 4(4):367–396, October 1974. CODEN ????

ISSN 0036-8539. URL https://www.jstor.org/stable/284547.

Werskey:1971:BSO

[Wer71] Paul Gary Werskey. British scientists and ‘outsider’ politics, 1931– 1945. Science Studies, 1(1):67–83, January 1971. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/370197.

Wolek:1974:PIC

[WG74] Francis W. Wolek and Belver C. Griffith. Policy and informal communications in applied science and technology. Science Studies, 4 (4):411–420, October 1974. CODEN ???? ISSN 0036-8539. URL

https://www.jstor.org/stable/284550.

Williams:1973:SPA

[Wil73] Roger Williams. Some political aspects of the Rothschild Affair. Science Studies, 3(1):31–46, January 1973. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/284463.

Wojick:1974:NRR

[Woj74] David Wojick. The norm of rationality or the rationality of norms. Science Studies, 4(2):193–195, April 1974. CODEN ???? ISSN 0036-

8539. URL https://www.jstor.org/stable/284476.

Woodward:1974:SGL

[Woo74] William R. Woodward. Scientific genius and loss of a parent. Science Studies, 4(3):265–277, July 1974. CODEN ???? ISSN 0036-8539.

URL https://www.jstor.org/stable/284525.

Weimer:1973:PNS

[WP73] Walter B. Weimer and David S. Palermo. Paradigms and normal science in psychology. Science Studies, 3(3):211–244, July 1973. CO-

DEN ???? ISSN 0036-8539. URL https://www.jstor.org/stable/
284494.

Weimer:1974:SSD

[WP74] Walter Weimer and David Palermo. Standards, scholarship, and debate: A rejoinder to Warren. Science Studies, 4(2):198–200, April 1974. CODEN ???? ISSN 0036-8539. URL https://www.jstor.

org/stable/284478.

Young:1971:EBI

[You71] Robert M. Young. Evolutionary biology and ideology: Then and now. Science Studies, 1(2):177–206, April 1971. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/370205.

Zinberg:1971:WGA

[Zin71] Dorothy Zinberg. The widening gap: Attitudes of first-year students and staff towards chemistry, science, careers and commitment. Science Studies, 1(3/4):287–313, October 1971. CODEN ???? ISSN

0036-8539. URL https://www.jstor.org/stable/370215.